# Dynamical Low-Rank Compression of Neural Networks with Robustness under Adversarial Attacks

### v1

@@ 4 A rank-adaptive, adversarial robustness increasing dynamical low-rank training scheme
Thus we adapt the two-step scheme of schotthöfer2024federateddynamicallowranktraining which ensures orthogonality of $U,V$. The method dynamically reduces or increases the rank of the factorized layers dynamically depending on the training dynamics and the complexity of the learning problem at hand. Consequently, the rank of each layer is no longer a hyper-parameter that needs fine-tuning, c.f. hu2021lora ; savostianova2023robustlowranktrainingapproximate , but is rather an interpretable measure for the inherent complexity required for each layer.
@@
To facilitate the discussion, we define ~$\widetilde{\mathcal{L}}=\mathcal{L}+\beta\mathcal{R}$ as the regularized loss function of the training process with regularization parameter $\beta>0$. To construct the method we consider the (stochastic) gradient descent-based update of a single weight matrix $W_{t+1}=W_{t+1}-\lambda\nabla_{W}\widetilde{\mathcal{L}}$ for minimizing $\widetilde{\mathcal{L}}$ with step size $\lambda>0$. The corresponding continuous time gradient flow reads $\dot{W}(t)=-\nabla_{W}\widetilde{\mathcal{L}}(W(t))$, which is a high-dimensional dynamical system with a steady state solution. We draw from established dynamical low-rank approximation (DLRA) methods, which were initially proposed for matrix-valued equations KochLubich07 . DLRA was recently extended to neural network training steffen2022low ; zangrando2023rank ; schotthöfer2024federateddynamicallowranktraining ; schotthöfer2024geolorageometricintegrationparameter ; kusch2025augmentedbackwardcorrectedprojectorsplitting ; Hnatiuk to formulate a consistent gradient flow evolution for the low-rank factors $U$, $S$, and $V$.
@@
The DLRA method constrains the trajectory of $W$ to the manifold $\mathcal{M}_{r}$, consisting of $n\times n$ matrices with rank $r$, by projecting the full dynamics $\dot{W}$ onto the local tangent space of $\mathcal{M}_{r}$ via an orthogonal projection, see Figure 3. The low-rank matrix is represented as $USV^{\top}\in\mathcal{M}_{r}$, where $U\in\mathbb{R}^{n\times r}$ and $V\in\mathbb{R}^{n\times r}$ have orthonormal columns and $S\in\mathbb{R}^{r\times r}$ is full-rank (but not necessarily diagonal). An explicit representation of the tangent space leads to equations for the factors $U$, $S$, and $V$ in (KochLubich07, , Proposition 2.1). However, following these equations requires a prohibitively small learning rate due to the curvature of the manifold lubich2014projector . Therefore, specialized integrators have been developed to accurately navigate the manifold with reasonable learning rates lubich2014projector ; ceruti2022unconventional ; ceruti2022rank .
@@
Below we summarize proposed method, called RobustDLRT, along with the changes induced by adding our robustness regularizer. A single iteration of DLRT is specified in Algorithm 1.
@@
Figure 3: Geometric interpretation of Algorithm 1. First, we compute the parametrization of the tangent plane $\mathcal{T}_{\mathcal{M}_{r}}$. Then we compute the projected gradient update with $\nabla_{\widehat{S}}\mathcal{L}$. Lastly, we retract the updated coefficients back onto the manifold $\mathcal{M}_{r}$. The regularizer $\mathcal{R}$ effectively changes the local curvature of $\mathcal{M}_{r}$.
@@
${\widehat{U}}&=\texttt{orth}([U^{t}\mid\nabla_{U}\mathcal{L}(U^{t}S^{t}V^{t,\top})])\in\mathbb{R}^{n\times 2r},\\ {\widehat{V}}&=\texttt{orth}([V^{t}\mid\nabla_{V}\mathcal{L}(U^{t}S^{t}V^{t,\top})])\in\mathbb{R}^{n\times 2r},\end{aligned}$ (6)
to double the rank of the low-rank representation and subsequently creates orthonormal bases ${\widehat{U}},{\widehat{V}}$. Since ${\mathcal{R}}(USV^{\top})={\mathcal{R}}(S)$, $\nabla_{U}{\mathcal{R}}(USV^{\top})=\nabla_{V}{\mathcal{R}}(USV^{\top})=0$; hence $\nabla_{U}\widetilde{\mathcal{L}}=\nabla_{U}\mathcal{L}$ and $\nabla_{V}\widetilde{\mathcal{L}}=\nabla_{V}\mathcal{L}$ are used in (6). The span on ${\widehat{U}}$ contains $U$, which is needed to ensure of the loss does not increase during augmentation, and a first-order approximation of the basis for $W^{t+1}$ using the exact gradient flow for $U$, see (schotthöfer2024federateddynamicallowranktraining, , Theorem 2) for details. Geometrically, the augmented space
@@
can be seen as subspace [^2] Technically the augmented space contains extra elements not in the tangent space, but this extra information only helps the approximation. of the tangent plane of $\mathcal{M}_{r}$ at $U^{t}S^{t}V^{t,\top}$, see Figure 3.
@@
We update the augmented coefficients ${\widehat{S}}$ via a Galerkin projection of the training dynamics onto the latent space $\mathcal{S}$. The augmented coefficients ${\widehat{S}}$ are updated by integrating the projected gradient flow $\dot{{\widehat{S}}}=-{\widehat{U}}^{\top}\nabla_{W}\widetilde{\mathcal{L}}{\widehat{V}}=-\nabla_{\widehat{S}}\widetilde{\mathcal{L}}$ using stochastic gradient descent or an other suitable optimizer for a number of $s_{*}$ local iterations, i.e.
${{\widehat{S}}_{s+1}}$ $={\widehat{S}}_{s}-\lambda\nabla_{\widehat{S}}\mathcal{L}-\beta\nabla_{\widehat{S}}\mathcal{R}({\widehat{S}}_{s}),$ (8)
Equation 8 is initialized with ${{\widehat{S}}_{0}}={\widehat{U}}^{\top}U^{t}S^{t}V^{t,\top}{\widehat{V}}\in\mathbb{R}^{2r\times 2r}$, and we set $\tilde{S}=\hat{S}_{s_{*}}$
@@
Finally, the augmented solution ${\widehat{U}}\tilde{S}{\widehat{V}}^{\top}$ is retracted back onto the manifold $\mathcal{M}_{r}$. The retraction can be computed efficiently by using a truncated SVD of $\tilde{S}$ that discards the smallest $r$ singular values. To enable rank adaptivity, the new rank $r_{1}$ instead of $r$ can be chosen by a variety of criteria, e.g., a singular value threshold $\left\lVert[\sigma_{r_{1}},\dots,\sigma_{2r}]\right\rVert_{2}<\vartheta$. Once a suitable rank is determined, the bases $U$ and $V$ are updated by discarding the basis vectors corresponding to the truncated singular values.
@@
The computational cost of the above described scheme is asymptotically the same as LoRA, since the reconstruction of the full weight matrix $W$ is never required. The orthonormalization accounts for $\mathcal{O}(nr^{2})$, the regularizer $\mathcal{R}$ for $\mathcal{O}(r^{3})$, and the SVD for $\mathcal{O}(r^{3})$ floating point operations. When using multiple coefficient update steps $s_{*}>1$, the amortized cost is indeed lower than that of LoRA, since only the gradient w.r.t ${\widehat{S}}$ is required in most updates.
@@
${\widehat{S}}\leftarrow{\widehat{U}}^{\top}USV^{\top}{\widehat{V}}$
@@ 3 $U,S,V\leftarrow$truncation$({\widehat{S}},{\widehat{U}},{\widehat{V}})$
9 def  truncation(${\widehat{S}}$: augmented coefficient, ${\widehat{U}}$: augmented basis, ${\widehat{V}}$: augmented co-basis ):
10       $P_{r_{1}},\Sigma_{r_{1}},Q_{r_{1}}\leftarrow$ truncated svd$(\widetilde{S})$ with threshold $\vartheta$ to new rank $r_{1}$
@@
        /* Coefficient update with diagonal $\Sigma_{r_{1}}$/
@@ 4.1 Extension to convolutional neural networks
The convolution layer map in 2D [^3] The notation extends to higher dimensional convolutions in a straight forward manner. CNNs translates a $W\times H$ image with $N_{I}$ in-features to $N_{O}$ out-features. Using tensors, this map is expressed as $Y=C*X$ where $X\in\mathbb{R}^{N_{I}\times W\times H}$, $Y\in\mathbb{R}^{N_{O}\times W\times H}$, and $C\in\mathbb{R}^{N_{O}\times N_{I}\times S_{W}\times S_{H}}$ is the convolutional kernel with a convolution window size $S_{W}\times S_{H}$. Neglecting the treatment of strides and padding, $C*X$ is given as a tensor contraction by
@@
where $s_{w}$ and $s_{h}$ range from $-S_{W}/2,\ldots,S_{W}/2$ and $-S_{H}/2,\ldots,S_{H}/2$ respectively, and $o=1,\dots,N_{O}$, $w=1,\dots,W$, and $h=1,\dots,H$.
@@
Aside from the prolongation (11a) and retraction (11c) from/to the low-rank latent space, the low-rank convolution map (11) equations features a convolution (11b) similar to (9) but in the reduced dimension low-rank latent space.
@@
The contractions in (9) and (11b) show that the output channels arise from a tensor contraction of the input channel and window modes; hence, both (9) and (11b) can be viewed as matrix-vector multiplication where $C$ is matricised on the output channel mode; i.e., $C\to\text{Mat}(C)\in\mathbb{R}^{N_{O}\times N_{I}S_{W}S_{H}}$ and $S\to\text{Mat}(S)\in\mathbb{R}^{r_{O}\times r_{I}S_{W}S_{H}}$. Therefore, we only regularize $\text{Mat}(S)$ with our robustness regularizer. Moreover, we assume $r_{O}\leq r_{I}S_{W}S_{H}$, which is almost always the case since $r_{O}$ and $r_{I}$ are comparable and $S_{W}S_{H}\gg 1$. Then we regularize convolutional layers by ${\mathcal{R}}(\text{Mat}(S)^{\top})$ so that $SS^{\top}$ is an $r_{O}\times r_{O}$ matrix, which is computationally efficient.
We remark, that the extension of Algorithm 1 to a tensor-valued layer with Tucker factorization only requires to change the truncation step; the SVD is replaced by a truncated Tucker decomposition of $S$. The Tucker bases $U_{i}$ can be augmented in parallel analogously to $U,V$ in the matrix case.
Table 2: UCM and Cifar10 benchmark. Clean and adversarial accuracy means and std. devs. of the baseline and regularized low-rank networks for different architectures. We report the low-rank results for unregulated, $\beta=0.0$, and the best performing $\beta$, given in Table 6. Algorithm 1 is able to match or surpass baseline adversarial accuracy values at compression rates of up to $95\%$ in most setups.
@@
VGG16 Baseline 0.0 94.40$\pm$0.72 86.71$\pm$1.90 76.40$\pm$2.84 54.96$\pm$2.99 89.58$\pm$2.99 85.05$\pm$3.40 77.77$\pm$1.61 37.25$\pm$3.66 23.05$\pm$3.01
DRLT 95.30 93.92$\pm$0.23 87.95$\pm$1.02 72.41$\pm$2.08 43.39$\pm$4.88 83.99$\pm$1.22 67.41$\pm$1.63 85.79$\pm$1.51 40.42$\pm$2.89 20.13$\pm$2.92
RobustDLRT 95.84 94.61$\pm$0.35 89.12$\pm$1.33 78.68$\pm$2.30 53.30$\pm$3.14 88.33$\pm$1.20 79.81$\pm$0.93 90.33$\pm$0.90 70.12$\pm$3.08 47.31$\pm$2.78
VGG11 Baseline 0.0 94.23$\pm$0.71 89.93$\pm$1.33 78.66$\pm$2.46 39.45$\pm$2.98 90.25$\pm$1.66 85.24$\pm$1.90 83.10$\pm$1.47 40.34$\pm$4.88 22.01$\pm$3.21
DRLT 94.89 93.70$\pm$0.71 86.58$\pm$1.22 67.55$\pm$2.16 28.92$\pm$2.65 83.90$\pm$1.36 63.41$\pm$1.39 87.15$\pm$1.18 40.17$\pm$4.96 14.18$\pm$3.78
RobustDLRT 94.59 93.57$\pm$0.84 87.90$\pm$0.91 72.96$\pm$1.55 32.85$\pm$2.46 86.77$\pm$0.76 74.31$\pm$1.50 88.00$\pm$1.13 60.97$\pm$4.18 28.56$\pm$3.64
ViT-32b Baseline 0.0 96.72$\pm$0.36 93.02$\pm$0.38 92.18$\pm$0.31 89.71$\pm$0.28 93.71$\pm$1.22 93.21$\pm$1.17 89.62$\pm$1.81 51.05$\pm$3.17 43.91$\pm$3.97
DRLT 86.7 96.38$\pm$0.60 91.21$\pm$0.44 82.10$\pm$0.32 62.45$\pm$0.41 86.67$\pm$1.05 79.81$\pm$0.81 80.48$\pm$1.82 41.52$\pm$3.24 35.91$\pm$3.76
RobustDLRT 87.9 96.41$\pm$0.67 92.57$\pm$0.34 85.67$\pm$0.41 69.94$\pm$0.42 91.03$\pm$0.86 84.19$\pm$1.39 87.33$\pm$1.81 46.39$\pm$2.75 40.76$\pm$3.88
@@
VGG16 Baseline 0.0 89.82$\pm$0.45 76.22$\pm$1.38 63.78$\pm$2.01 34.97$\pm$2.54 78.60$\pm$1.12 73.54$\pm$1.55 71.51$\pm$1.31 37.36$\pm$2.60 16.12$\pm$2.12
DRLT 94.37 89.23$\pm$0.62 74.07$\pm$1.23 59.55$\pm$1.79 28.74$\pm$2.21 72.51$\pm$1.04 66.21$\pm$1.41 79.56$\pm$1.15 59.88$\pm$2.26 38.98$\pm$1.94
RobustDLRT 94.18 89.49$\pm$0.58 76.04$\pm$1.18 62.08$\pm$1.69 32.77$\pm$2.04 75.53$\pm$0.98 69.93$\pm$1.22 87.62$\pm$1.07 84.80$\pm$2.01 81.26$\pm$2.15
VGG11 Baseline 0.0 88.34$\pm$0.49 75.89$\pm$1.42 64.21$\pm$1.96 31.76$\pm$2.45 74.96$\pm$1.09 68.59$\pm$1.63 74.77$\pm$1.26 40.88$\pm$2.58 08.95$\pm$1.98
DRLT 95.13 88.13$\pm$0.56 72.02$\pm$1.34 55.83$\pm$1.92 21.59$\pm$2.16 66.98$\pm$1.05 58.57$\pm$1.55 79.42$\pm$1.08 47.95$\pm$2.18 22.92$\pm$1.77
RobustDLRT 94.67 87.97$\pm$0.52 76.04$\pm$1.26 63.82$\pm$1.83 30.77$\pm$2.30 71.06$\pm$1.00 65.63$\pm$1.38 84.93$\pm$1.10 78.35$\pm$1.89 65.93$\pm$2.04
ViT-32b Baseline 0.0 95.42$\pm$0.35 79.94$\pm$0.95 63.66$\pm$1.62 32.09$\pm$2.05 84.65$\pm$0.88 77.20$\pm$1.04 52.17$\pm$1.49 16.03$\pm$2.34 13.29$\pm$2.01
DRLT 73.42 95.39$\pm$0.41 79.50$\pm$0.91 61.62$\pm$1.48 30.32$\pm$1.94 83.33$\pm$0.80 76.16$\pm$0.95 58.32$\pm$1.44 17.43$\pm$2.28 14.49$\pm$1.92
RobustDLRT 75.21 94.66$\pm$0.38 82.03$\pm$0.88 69.29$\pm$1.43 38.05$\pm$1.99 87.97$\pm$0.75 83.03$\pm$0.91 74.49$\pm$1.32 27.80$\pm$2.11 18.34$\pm$1.87

### v2

@@ 5 A rank-adaptive, adversarial robustness increasing dynamical low-rank training scheme
Thus we adapt the two-step scheme of schotthöfer2024federateddynamicallowranktraining which ensures orthogonality of $U,V$. The method dynamically reduces or increases the rank of the factorized layers depending on the training dynamics and the complexity of the learning problem at hand. Consequently, the rank of each layer is no longer a hyper-parameter that needs fine-tuning, c.f. hu2021lora ; savostianova2023robustlowranktrainingapproximate , but is rather an interpretable measure for the inherent complexity required for each layer.
@@
To facilitate the discussion, we define $\widetilde{\mathcal{L}}=\mathcal{L}+\beta\mathcal{R}$ as the regularized loss function of the training process with regularization parameter $\beta>0$. To construct the method we consider the (stochastic) gradient descent-based update of a single weight matrix $W_{t+1}=W_{t+1}-\lambda\nabla_{W}\widetilde{\mathcal{L}}$ for minimizing $\widetilde{\mathcal{L}}$ with step size $\lambda>0$. The corresponding continuous time gradient flow reads $\dot{W}(t)=-\nabla_{W}\widetilde{\mathcal{L}}(W(t))$, which is a high-dimensional dynamical system with a steady state solution. We draw from established dynamical low-rank approximation (DLRA) methods, which were initially proposed for matrix-valued equations KochLubich07 . DLRA was recently extended to neural network training steffen2022low ; zangrando2023rank ; schotthöfer2024federateddynamicallowranktraining ; schotthöfer2024geolorageometricintegrationparameter ; kusch2025augmentedbackwardcorrectedprojectorsplitting ; Hnatiuk to formulate a consistent gradient flow evolution for the low-rank factors $U$, $S$, and $V$.
@@
The DLRA method constrains the trajectory of $W$ to the manifold $\mathcal{M}_{r}$, consisting of $n\times n$ matrices with rank $r$, by projecting the full dynamics $\dot{W}$ onto the local tangent space of $\mathcal{M}_{r}$ via an orthogonal projection, see Figure˜3. The low-rank matrix is represented as $USV^{\top}\in\mathcal{M}_{r}$, where $U\in\mathbb{R}^{n\times r}$ and $V\in\mathbb{R}^{n\times r}$ have orthonormal columns and $S\in\mathbb{R}^{r\times r}$ is full-rank (but not necessarily diagonal). An explicit representation of the tangent space leads to equations for the factors $U$, $S$, and $V$ in (KochLubich07, , Proposition 2.1). However, following these equations requires a prohibitively small learning rate due to the curvature of the manifold lubich2014projector . Therefore, specialized integrators have been developed to accurately navigate the manifold with reasonable learning rates lubich2014projector ; ceruti2022unconventional ; ceruti2022rank .
@@
Below we summarize the proposed method, called RobustDLRT, along with the changes induced by adding our robustness regularizer. A single iteration of RobustDLRT is specified in Algorithm˜1.
@@
Figure 3: Geometric interpretation of Algorithm˜1. First, we compute the parametrization of the tangent plane $\mathcal{T}_{\mathcal{M}_{r}}$. Then we compute the projected gradient update with $\nabla_{\widehat{S}}\mathcal{L}$. Lastly, we retract the updated coefficients back onto the manifold $\mathcal{M}_{r}$. The regularizer $\mathcal{R}$ steers training to regions of $\mathcal{M}_{r}$ with lower curvature.
@@
$\begin{aligned} {\widehat{U}}&=\texttt{orth}([U^{t}\mid\nabla_{U}\mathcal{L}(U^{t}S^{t}V^{t,\top})])\in\mathbb{R}^{n\times 2r},\\ {\widehat{V}}&=\texttt{orth}([V^{t}\mid\nabla_{V}\mathcal{L}(U^{t}S^{t}V^{t,\top})])\in\mathbb{R}^{n\times 2r},\end{aligned}$ (6)
to double the rank of the low-rank representation and subsequently creates orthonormal bases $^{\widehat{U}},{\widehat{V}}$. Since ${\mathcal{R}}(USV^{\top})={\mathcal{R}}(S)$, $\nabla_{U}{\mathcal{R}}(USV^{\top})=\nabla_{V}{\mathcal{R}}(USV^{\top})=0$; hence $\nabla_{U}\widetilde{\mathcal{L}}=\nabla_{U}\mathcal{L}$ and $\nabla_{V}\widetilde{\mathcal{L}}=\nabla_{V}\mathcal{L}$ are used in (6). The span of $^{\widehat{U}}$ contains $U^{t}$, which is needed to ensure of the loss does not increase during augmentation, and a first-order approximation of $\mathrm{span}(U^{t+1})$ using the exact gradient flow for $U$, see (schotthöfer2024federateddynamicallowranktraining, , Theorem 2) for details. Geometrically, the latent space
@@
can be seen as subspace [^3] Technically the latent space contains extra elements not in the tangent space, but the extra information only helps the approximation. of the tangent plane of $\mathcal{M}_{r}$ at $U^{t}S^{t}V^{t,\top}$, see Figure˜3.
@@
We update the latent coefficients $^{\widehat{S}}$ via a Galerkin projection of the training dynamics onto the latent space $\mathcal{S}$. The latent coefficients $^{\widehat{S}}$ are updated by integrating the projected gradient flow $\dot{{\widehat{S}}}=-{\widehat{U}}^{\top}\nabla_{W}\widetilde{\mathcal{L}}{\widehat{V}}=-\nabla_{\widehat{S}}\widetilde{\mathcal{L}}$ using stochastic gradient descent or an other suitable optimizer for a number of $s_{*}$ local iterations, i.e.
${{\widehat{S}}_{s+1}}$ $={\widehat{S}}_{s}-\lambda\nabla_{\widehat{S}}\mathcal{L}-\beta\nabla_{\widehat{S}}\mathcal{R}({\widehat{S}}_{s}),\quad s=0,\ldots,s_{*}-1.$ (8)
Equation˜8 is initialized with ${{\widehat{S}}_{0}}={\widehat{U}}^{\top}U^{t}S^{t}V^{t,\top}{\widehat{V}}\in\mathbb{R}^{2r\times 2r}$, and we set $\tilde{S}=\hat{S}_{s_{*}}$
@@
Finally, the latent solution $^{\widehat{U}}\tilde{S}{\widehat{V}}^{\top}$ is retracted back onto the manifold $\mathcal{M}_{r}$. The retraction can be computed efficiently by using a truncated SVD of $\tilde{S}$ that discards the smallest $r$ singular values. To enable rank adaptivity, the new rank $r_{1}$ instead of $r$ can be chosen by a variety of criteria, e.g., a singular value threshold $\left\lVert[\varsigma_{r_{1}},\dots,\varsigma_{2r}]\right\rVert_{2}<\vartheta$. Once a suitable rank is determined, the bases $U$ and $V$ are updated by discarding the basis vectors corresponding to the truncated singular values.
Remark 3.
[…]
@@
The computational cost of RobustDLRT is asymptotically the same as LoRA, since the reconstruction of the full weight matrix $W$ is never required. The orthonormalization accounts for $\mathcal{O}(nr^{2})$, the regularizer $\mathcal{R}$ for $\mathcal{O}(r^{3})$, and the SVD for $\mathcal{O}(r^{3})$ floating point operations. When using multiple coefficient update steps $s_{*}>1$, the amortized cost is indeed lower than that of LoRA, since only the gradient with respect to $^{\widehat{S}}$ is required in most updates. While the regularizer may be applied to full-rank baseline models, its $\mathcal{O}(n^{3})$ computational scaling significantly increases training costs.
@@
$^{\widehat{S}}\leftarrow{\widehat{U}}^{\top}USV^{\top}{\widehat{V}}$
@@
9 def truncation($^{\widehat{S}}$: augmented coefficient, $^{\widehat{U}}$: augmented basis, $^{\widehat{V}}$: augmented co-basis ):
10 $P_{r_{1}},\Sigma_{r_{1}},Q_{r_{1}}\leftarrow$ truncated svd(S~)($\widetilde${S}) with threshold $\vartheta$ to new rank $r_{1}$
@@
 /* Coefficient update with diagonal $\Sigma_{r_{1}}$ */
@@ 5.1 Extension to convolutional neural networks
The convolution layer map in 2D CNNs translates a $W\times H$ image with $N_{I}$ in-features to $N_{O}$ out-features. Using tensors, this map is expressed as Y=CXY=C*X where $X\in\mathbb{R}^{N_{I}\times W\times H}$, $Y\in\mathbb{R}^{N_{O}\times W\times H}$, and $C\in\mathbb{R}^{N_{O}\times N_{I}\times S_{W}\times S_{H}}$ is the convolutional kernel with a convolution window size $S_{W}\times S_{H}$. Neglecting the treatment of strides and padding, CXC*X is given as a tensor contraction by
@@
where $s_{w}$ and $s_{h}$ range from $S_{W}/2,\ldots,S_{W}/2$ and $S_{H}/2,\ldots,S_{H}/2$ respectively, and $o=1,\dots,N_{O}$, $w=1,\dots,W$, and $h=1,\dots,H$.
@@
Aside from the prolongation (11a) and retraction (11c) from/to the low-rank latent space, the low-rank convolution map (11) features a convolution (11b) similar to (9) but in the reduced dimension low-rank latent space.
@@
The contractions in (9) and (11b) show that the output channels arise from a tensor contraction of the input channel and window modes; hence, both (9) and (11b) can be viewed as matrix-vector multiplications where $C$ is matricised on the output channel mode; i.e., $C\to\text{Mat}(C)\in\mathbb{R}^{N_{O}\times N_{I}S_{W}S_{H}}$ and $S\to\text{Mat}(S)\in\mathbb{R}^{r_{O}\times r_{I}S_{W}S_{H}}$. Therefore, we only regularize $\text{Mat}(S)$ with our robustness regularizer. Moreover, we assume $r_{O}\leq r_{I}S_{W}S_{H}$, which is almost always the case since $r_{O}$ and $r_{I}$ are comparable and $S_{W}S_{H}\gg 1$. Then we regularize convolutional layers by ${\mathcal{R}}(\text{Mat}(S)^{\top})$ so that $SS^{\top}$ is an $r_{O}\times r_{O}$ matrix, which is computationally efficient.
We remark that the extension of Algorithm˜1 to a tensor-valued layer with Tucker factorization only requires to change the truncation step; the SVD is replaced by a truncated Tucker decomposition of $S$. The Tucker bases $U_{i}$ can be augmented in parallel similarly to the matrix case.
Table 2: UCM and Cifar10 benchmark. Clean and adversarial accuracy means and std. devs. of the baseline and regularized low-rank networks for different architectures. We report the low-rank results for unregulated, $\beta=0.0$, and the best performing $\beta$, given in Table˜9. Algorithm˜1 (RobustDLRT)  is able to match or surpass baseline adversarial accuracy values at compression rates of up to $95\%$ in most setups. All runs where RobustDLRT surpasses the uncompressed baseline are highlighted.
@@
VGG16
[…]
VGG11
[…]
ViT-16b
[…]
@@
VGG16
[…]
VGG11
[…]
ViT-16b
[…]